A large information technology (IT) company sells a software product designed to help protect an organization’s data. The company’s market research has determined that hospitals and other medical organizations in its market area have a 1.3 percent chance of losing some of a patient’s data over a five-year period. With the IT company’s software, the chances of patient data loss for a five-year period are reduced to 0.12 percent. Further, the market research has determined that, when patient data loss occurs, there is a 15 percent chance that the organization will lose the patient’s business and a 0.75 percent chance that a lawsuit will result. The average cost of losing a patient’s business is $600 in promotion expenses and other start-up costs to acquire a new patient. The average cost of a lawsuit is $100,000. Given these research findings, calculate the value of this data-protection benefit of the IT company’s software product to a medical organization that keeps data records on 20,000 patients.
Answer #1

Answer:

Estimation of IT companies information insurance programming = Net reserve funds on information loss

= cost of information loss without programming acquisition - cost of information loss with programming acquisition.

1. COST IN SOFTWARE NOT ACQUIRED

No of patients = 20000

Data loss = 20000*1.3% = 260

Cost of loss due to patient's business = 260*15%*600 = $23400

Cost of filing law suit = 260*0.75%*100000 = $195000

Total cost if software not acquired = 23400+195000 = $218400

2. COST IF SOFTWARE ACQUIRED

No of patients = 20000

Data loss = 20000*0.12% = 24

Cost of loss due to patients' business = 24*15%*600 = $2160

Cost of filing law suit = 24*0.75%100000 =$18000

Total cost if software acquired = 2160+18000 = $20160

-> Values of IT companies data protection software = Net savings data loss

= Cost of data loss without software acquisition - Cost of data loss with software acquisition

= 218400 - 20160 = $198240